OpenStudy (anonymous):

classify the equation as an identity, conditional equation, or contradiction. if it is a conditional equation, estimate the smallest positive solution, cosxsinx=cosx+sinx

OpenStudy (math&ing001):

The points where the 2 functions meet are the solutions to your equation.

OpenStudy (anonymous):

how do you find those 2 points out though?

OpenStudy (anonymous):

nevermind i figured it out. thanks! can you also help me with this question:

OpenStudy (math&ing001):

There are more than 2 points as solutions to your equation, in fact there's an infinity of solutions. But you're only asked to give the first positive one.

OpenStudy (anonymous):

find the expression that is equal to \[\tan ^{2}w+3\]

OpenStudy (math&ing001):

You could write it as: \[\tan ^{2}w+3=\sec ^{2}w+2\]

OpenStudy (anonymous):

okay what about sinw/secw+1 + sinw/secw-1

OpenStudy (anonymous):

its the same type of question.

OpenStudy (math&ing001):

sin(w)/sec(w)=sin(w)*cos(w)=(1/2)sin(2w) Hope that helps.

OpenStudy (anonymous):

the choices are: cscw+cotw^2 tan w 2cotw 1-csc^2w 1-cot^2w

OpenStudy (math&ing001):

Did you mean secw+1 as sec(w+1) ?

OpenStudy (anonymous):

no there are no parentheses

OpenStudy (math&ing001):

sinw/secw+1 + sinw/secw-1 Can you write it better or draw it ?

OpenStudy (anonymous):